OpenStudy (mendicant_bias):

Taking a fairly ugly final IVP dealing with Dirac-Delta before moving onto my last subject to cover, Systems of Diff. Eqn's. Posted below in a moment.

OpenStudy (mendicant_bias):

Oh! Now I know why the argument was just sin(t)! It's because it makes a full round and it doesn't actually affect the outcome of the sine function on that last problem! sin(t-2pi)=sin(t)!

OpenStudy (mendicant_bias):

(Lol now onto this problem)

OpenStudy (sidsiddhartha):

yeah lol god!!

OpenStudy (mendicant_bias):

\[y''+4y'+13y=\delta(t-\pi)+\delta(t-3\pi), \ \ \ y(0)=1, \ y'(0)=0\]

OpenStudy (mendicant_bias):

(That doesn't look as bad as I thought but w/e, might get bad soon)

OpenStudy (mendicant_bias):

\[[s^2Y(s)-sy(0)-y'(0)]+4[sY(s)-y(0)]+13Y(s)=e^{-\pi s}+e^{-3\pi s}\]

OpenStudy (anonymous):

hmmmmmm............345678900987653673559

OpenStudy (mendicant_bias):

\[s^2Y(s)-s+4sY(s)-1+13Y(s)=e^{-\pi s}+e^{-3\pi s}\] \[Y(s)[s^2+4s+13]=e^{-\pi s}+e^{-3\pi s}+1+s\]

OpenStudy (mendicant_bias):

\[Y(s)=\frac{e^{-\pi s}+e^{-3\pi s}+1+s}{s^{2}+4s+13}\]

OpenStudy (mendicant_bias):

(So much for "doesn't look too bad")

OpenStudy (sidsiddhartha):

hey wait

OpenStudy (mendicant_bias):

Alright, now, splitting up the terms sounds sensible, but I have no idea what to do with that denominator, doesn't appear to be doable with partial fractions. Even if I don't solve this to the end, I just need to figure out a general idea. (And if you gotta go, I understand, getting late there, heh. Don't mean to keep you)

OpenStudy (mendicant_bias):

Mistake?

OpenStudy (sidsiddhartha):

no it fine go ahead

OpenStudy (sidsiddhartha):

*its

OpenStudy (mendicant_bias):

Alright, now not sure what to do

OpenStudy (mendicant_bias):

brb (>2 mins, short time)

OpenStudy (sidsiddhartha):

man this one is going to be (*dont know what to say*) lol

OpenStudy (sidsiddhartha):

\[s^2+4s+13=(s+2)^2+9\]

OpenStudy (sidsiddhartha):

that'll help

OpenStudy (mendicant_bias):

Alright, back

OpenStudy (sidsiddhartha):

\[\frac{ e^{-\pi s} }{ (s+2)^2+3^2 }+\frac{ e^{-3 \pi s} }{ (s+2)^2+3^2 }+\frac{ 1 }{ (s+2)^2+3^3 }+\frac{ s }{ (s+2)^2+3^3 }\]

OpenStudy (mendicant_bias):

Oh, okay, now evaluating {L^{-1}}

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{e^{-\pi s}}{(s+2)^2+3^2}\right\}=\](Using Heaviside and shifting) Well, actually, no, I'm not so sure how to evaluate this, I'm not sure if it applies to both inverse laplace transforms, but don't all instances of s have to be paired with the same subtraction/addition?

OpenStudy (sidsiddhartha):

yes thats the idea

OpenStudy (mendicant_bias):

So do I need a\[e^{-\pi(s+2)}\]in the denominator?

OpenStudy (sidsiddhartha):

no u dont because it does'nt affect exponentials

OpenStudy (mendicant_bias):

Alright, great! So then I can just apply the shift as a multiple of e^{-2t} or something like that on the t-axis, or some kind of shifting again? (Gonna have to look it up, one moment)

OpenStudy (mendicant_bias):

Got it, First Translation Theorem, \[\mathcal{L}\left\{e^{at}f(t)\right\}=F(s-a)\]

OpenStudy (sidsiddhartha):

yes for the first term\[e^{-2(t-\pi)}\sin3(t-\pi).u(t-\pi)\] that looks ok?

OpenStudy (mendicant_bias):

heh, lol, does that work out by complete chance? One sec, working it myself.

OpenStudy (mendicant_bias):

\[e^{-2(t-\pi)}\sin(3(t-2\pi))\mathcal{U}(t-2\pi)\]Somewhere we're missing a fraction, that whole thing is supposed to be multiplied by 1/3rd

OpenStudy (mendicant_bias):

Whoops, pi, not 2pi

OpenStudy (mendicant_bias):

I'm just wondering where we're missing that 1/3rd that the whole thing is multiplied by.

OpenStudy (sidsiddhartha):

oooooohhhhhhhhhh yes

OpenStudy (sidsiddhartha):

i forgot

OpenStudy (sidsiddhartha):

\[L[\sin3t]=\frac{ 3 }{ s^2+9 }\]

OpenStudy (sidsiddhartha):

look we did'nt multiplied that 1/3 to make it sib(3t)

OpenStudy (mendicant_bias):

Oh! I see!

OpenStudy (mendicant_bias):

Alright, well, cool, I'll have time to take time on my test/check for algebra mistakes, for now I'm just covering concepts; I'm going to move onto my final subject, systems of Diff Eqns. never done any, know nothing about it, just a quick read and working through practice problems.

OpenStudy (sidsiddhartha):

yes those are not that hard

OpenStudy (mendicant_bias):

Alright, cool. Getting close to the end. Just let me know if you have to go, must be getting tired over there.

OpenStudy (sidsiddhartha):

yes sorry i've to go :( its now 2 am and gotta sleep :) bye :)